how do i solve the following equation for x: 2 - 4(3x + 1) = 8 - 7x
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

2-4(3X+1)=8-7X
2-12X-4=8-7X
-12X+7X=8-2+4
-5X=10
X=10/-5
X=-2 ANSWER.
PROOF
2-4(3*-2+1)=8-7*-2
2-4(-6+1)=8+14
2-4(-5)=22
2+20=22
22=22

Answer by stanbon(75887)

solve the following equation for x: 2 - 4(3x + 1) = 8 - 7x
2 -12x -4 = 8 - 7x
-12x - 2 = 8 - 7x
5x=-10
x=-2
===========
Cheers,
Stan H.